If $P are invested at an interest rate r compounded annually, at the end of 2 years the amount will be A=P(1+r). At what interest rate will $1000 increase to $1440 in 2 years ? A=$1440 and P=$1000.

Answer by dfrazzetto(283)

According to the stated formula (which does not look how annually compounded interest works FYI):
A=P(1+r)
1440 = 1000(1+r)
1440 = 1000 + 1000r
1000r = 440
r= 440/1000=.44 = 44%